The probability density function of a random variable X is f(x) = C|x|. Find C, and the value of x° such that Fx(x°) = 3/4, where F is the CDF.


1
Expert's answer
2021-05-07T10:18:22-0400

The pdf is-

"f(x)=Cx, 0\le x\le 2"

"0 , \text{ othervise }"


As we know that for pdfs-


"\int_{-\infty}^{\infty}f(x)dx=1"


"\int_0^2Cxdx=1"


"\dfrac{Cx^2}{2}|_0^2=1"

"\dfrac{4C}{2}=1\Rightarrow C=\dfrac{1}{2}"



Also CDF is calculating by integrating the PDf-


"F_X(x^o)=\int_{-\infty}^{x^o}\dfrac{x}{2}dx"



"F_X(x^o)=\int_{0}^{x^o}\dfrac{x}{2}dx"



"\dfrac{3}{4}=\dfrac{x^2}{4}|_0^{x^o}\Rightarrow \dfrac{(x^o)^2}{4}=\dfrac{3}{4}\Rightarrow x^o=\sqrt{3}"